The length of a rectangular parking lot is 4 meters less than twice its width, and the perimeter is 412 meters, find the length of the parking lot.
Answer by algebrahouse.com(1659)

x = width
x - 4 = length

Perimeter of a rectangle = 2(width) + 2(length)

2x + 2(x - 4) = 412 {substituted into perimeter formula}
2x + 2x - 8 = 412 {used distributive property}
4x - 8 = 412 {combined like terms}
4x = 420 {added 8 to each side}
x = 105 {divided each side by 4}
x - 4 = 101 {substituted 130, in for x, into x - 4}

width = 105 m
length = 101 m

For more help from me, visit: www.algebrahouse.com